One leg of a right triangle is 5 cm and the hypotenuse measures 5 square root of 2 cm. Find the length of the other leg

Answer by ewatrrr(24785)

Isosceles Right Triangle:
(5√2)^2 = 2*5^2 = 5^2 + 5^2
.......
the length of the other leg: 5cm